A piece of copper 700g is heated to100°c and then quickly transferred to a copper calorimeter of mass 20g containing 500g of liquid of unknown specific heat capacity of liquid calculate the specific heat capacity of copper 3900j/kg




1
Expert's answer
2022-04-19T10:00:59-0400

"c_1m_1(100-50)=c_1m_2(50-20)+cm_3(50-20),"

"c=2500~\frac J{kg\cdot °C}."